OpenStudy (anonymous):

Write the equation of the tangent line to the graph of f at the given point / x-value f(x)=x^2+2x+1 @ (-3,4) Need someone to do the process because I don't understand how to solve it.

OpenStudy (mathmale):

Would you please find the derivative of your function f(x)? That derivative represents the slope of the tangent line to the curve.

OpenStudy (anonymous):

f'(x)=2x+2

OpenStudy (mathmale):

Cool. So, f'(x) = 2x + 2. Now please let x = -3. (Why?)

OpenStudy (anonymous):

umm to find the slope....

OpenStudy (mathmale):

Yes. But why use x=-3?

OpenStudy (anonymous):

i'm not sure

OpenStudy (mathmale):

I wanted you to make that connection. You're given the point (-3,4) and want to find the equation of the tangent line to the graph at that point. That's where the -3 comes from. You're looking to write the equation of the tangent line to the graph at (-3,4), right?

OpenStudy (anonymous):

yes

OpenStudy (mathmale):

Cool. so, if the derivative f'(x) is 2x+2, and x=-3, evaluate f'(-3).

OpenStudy (anonymous):

f'(-3)=-4

OpenStudy (mathmale):

That's it. What's the significance of this result?

OpenStudy (anonymous):

I suppose that -4 is the slope therefore if we put it in slope format y-y=m(x-x) the answer will be y-4=-4(x+3)

OpenStudy (mathmale):

That's right! -4 is the slope. Perfect. Please simplify your result. Good going!!

OpenStudy (anonymous):

Thank you

OpenStudy (mathmale):

Simplify y-4=-4(x+3): y=4-4(x+3). Care to take the simplification one step further?

OpenStudy (anonymous):

no this is fine, thank you